OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10? 605 660 748 816

OpenStudy (anonymous):

Is there any one smart enough to help me solve this problem??

OpenStudy (jack1):

\[a _{n} = a _{1} + (n-1)d\] arithmetic series equation

OpenStudy (jack1):

an = nth term a1 = first term d = difference between successive terms n = which term you are up to

OpenStudy (anonymous):

ok

OpenStudy (jack1):

so d = -10 the 12th term = 13 = a12 n = 12 so find a1...@dsolorzano

OpenStudy (anonymous):

1 sec

OpenStudy (jack1):

then plug that value into the following equation: sum = [n (a1 + an) ]/2

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

??

OpenStudy (jack1):

\[a _{n}=a _{1}+(n−1)d, where: a _{12} = 13\] \[a _{1} = ???\] \[(n-1)(d) = (12-1)\times (-10)\]

OpenStudy (anonymous):

(n-1)(d)=-110

OpenStudy (jack1):

\[a _{12} = a _{1}+(n−1)d\]

OpenStudy (jack1):

\[13 = a _{1}+ (-110)\]

OpenStudy (anonymous):

123=a1?

OpenStudy (jack1):

yep. now sub that into sum = [n (a1 + an) ]/2

OpenStudy (anonymous):

123={n(a1+an}/2

OpenStudy (jack1):

\[\sum of series = \frac{ n(a _{1}+a _{n}) }{ 2 }\]

OpenStudy (jack1):

remember 123 is a1 not sum of series

OpenStudy (anonymous):

ok

OpenStudy (jack1):

and an = 13 and n =12

OpenStudy (anonymous):

12(123+13)/2

OpenStudy (jack1):

yep, equals...?

OpenStudy (anonymous):

816

OpenStudy (jack1):

sweet as man

OpenStudy (anonymous):

Thanks for your help

OpenStudy (jack1):

so just remember those 2 equations: an=a1+(n−1)d for the value of term n (in this case the 12th term) and sum = n(an+a1)/2

OpenStudy (jack1):

welcome dude